OpenStudy (clamin):

PLEASE HELP!! f(x)=-2(x-2)^2+4 (graph)

OpenStudy (anonymous):

Can you tell me what each of the transformations are to get f(x)=x^2 to f(x)=-2(x-2)^2+4

OpenStudy (clamin):

-b/2a is the equation

OpenStudy (anonymous):

That's not an answer to my question

OpenStudy (triciaal):

for the general equation ax^2 + bx + c = 0 in vertex form f(x) = a(x-h)^2 + k h = -b/2a and k = b^2/4a^2 transformations -a means reflection across the x-axis -h means moves the graph right + h moves the graph left -k moves the graph down and + k moves the graph up

OpenStudy (triciaal):

a means expand or shrink

OpenStudy (triciaal):

to get f(x)=x^2 to f(x)=-2(x-2)^2+4 a = 2 twice as big , - =reflect across x-axis x-2 = 2 places right +4 = 4 places up